how many coins makes 6 dollars?
Answer by checkley77(12844)

.50*12=$6.00 12 half dollars.
.25*24=$6.00 24 quarters.
.10*60=$6.00 60 dimes.
.05*120=$6.00 120 nickles.
.01*600=$6.00 600 pennies.
Combinations of these coins can also make $6.00 but there are too many to list here in a table.